How many particles are present in a closed container if the energy it contains is 95577.88J, and the diatomic oxygen gas is moving at a velocity of 51.28m/s?


1
Expert's answer
2021-12-13T10:34:20-0500

Solution:


Energy and number of particle relation is given by as

"E=\frac{nM_Av^2}{2N_A}"

"n=\frac{2E N_A}{M_Av^2}"


"n=\frac{2\times95577.88\times 6.023\times 10^{23}}{32\times(51.28)^2}"


"n=13.68\times10^{23}"